# Document Valuation in LLM Summaries: A Cluster Shapley Approach

### v2

@@ Abstract
Large Language Models (LLMs) are increasingly used in systems that retrieve and summarize content from multiple sources, such as search engines and AI assistants. While these models enhance user experience by generating coherent summaries, they obscure the contributions of original content creators, raising concerns about credit attribution and compensation. We address the challenge of valuing individual documents used in LLM-generated summaries. We propose using Shapley values, a game-theoretic method that allocates credit based on each document’s marginal contribution. Although theoretically appealing, Shapley values are expensive to compute at scale. We therefore propose Cluster Shapley, an efficient approximation algorithm that leverages semantic similarity between documents. By clustering documents using LLM-based embeddings and computing Shapley values at the cluster level, our method significantly reduces computation while maintaining attribution quality. We demonstrate our approach to a summarization task using Amazon product reviews. Cluster Shapley significantly reduces computational complexity while maintaining high accuracy, outperforming baseline methods such as Monte Carlo sampling and Kernel SHAP by offering a better efficient frontier. Our approach is agnostic to the exact LLM used, the summarization process used, and the evaluation procedure, which makes it broadly applicable to a variety of summarization settings.
Keywords: LLMs, Shapley Value, Digital Marketing, Search System, Retrieval Augmented Generation, Reviews.
@@ 1 Introduction
The advent of Large Language Models (LLMs) has revolutionized how users search, process, and consume information. Today’s LLM-based search and summarization engines combine the strengths of LLMs with those of traditional search engines. First, unlike traditional search engines that return a list of links in response to a query, LLM-powered search provides a concise summary with references to relevant documents/websites. This frees users from the cognitive load of manually navigating and aggregating information from multiple sources (searchgpt). Second, unlike regular LLMs that rely solely on static training data, LLM-based search engines augment their generative process with real-time retrieval and source grounding using a framework known as Retrieval-Augmented Generation (RAG) (fan2024survey). In the RAG framework, the system first retrieves a set of documents/articles that are most relevant to the query and then uses them as context in the generative process. This ensures that the LLM’s responses are grounded in up-to-date, relevant content and do not suffer from limitations such as non-factual hallucinations and outdated knowledge that plague off-the-shelf LLMs. Thus, LLM-based search and summarization platforms combine the generative aspects of LLMs with the retrieval aspects of search engines by augmenting generation with context/information from documents most relevant to a query.
@@
Over the last few years, all major search engines have adopted these developments and integrated LLMs into their search infrastructures, e.g., Microsoft’s Bing AI (microsoft2025copilotsearch) and Google’s AI Overview (google2025ai-overviews). Both these tools provide a summary response to a user’s search query with references or links to relevant websites for informational queries. [^1] Search queries are typically categorized into three groups – (1) informational, where the user is seeking information on a topic, (2) navigational, where the user is simply using the search engine to navigate to a website, and (3) transactional, where the user seeks to perform some transaction, e.g., a purchase. LLM-based summaries are mainly shown for informational queries. Further, new LLM-based search engines such as OpenAI’s ChatGPT Search (searchgpt) and Perplexity AI (perplexityai2025) have emerged and grown rapidly in the last few years. The impact of this transformation goes beyond web search and extends to Q&A and informational websites (e.g., Reddit) and e-commerce websites (e.g., Amazon, Best Buy). For example, Amazon now shows an LLM-based summary of all the reviews for a product on the main product page. Customers can further access the source reviews the summary is built on, if needed; see Figure 1 for an example. In addition to broad overviews of the reviews, consumers can query Amazon’s AI-based search tool Rufus for specific information related to the product, and it provides a response based on the information in the reviews/product page (amazon_gen_ai_reviews; amazon_rufus). Examples of all these applications are illustrated in Web Appendix §A.
@@
The main advantage of LLM-based search and summarization tools is that they simplify the search process for users, who no longer have to click on multiple links and aggregate information from individual sources. This can increase customer satisfaction and lead to higher platform usage (xu2023chatgpt). (zhu2025generative) empirically shows that LLM-based search and summarization can significantly boost consumer purchases on an e-commerce platform through better articulation of consumer needs and product attributes. Furthermore, these approaches keep customers on the platform’s own website or interface; that is, consumers don’t need to leave the search engine and visit other websites to gather information. This, in turn, can lead to increased time spent on the platform, resulting in more eyeballs and potentially higher revenues (Goodwin2024). For these reasons, the rise of LLM-based search and summarization has generally been seen as beneficial to both platforms and consumers, at least in the short term.
@@
However, the rise of LLM-based automated summarization tools has led to significant concerns among content creators and information providers. In a traditional search model, the source documents (and their corresponding producers) stand to gain traffic and reputation by being shown in response to search queries in organic listings. [^2] Both higher traffic and better reputation can usually be monetized through advertising and/or subscription revenues. In contrast, LLM-powered search delivers a synthesized answer directly in the search interface, allowing users to obtain information without visiting the source websites. For example, after the introduction of Google’s AI overview, website owners have expressed reservations over the loss of traffic (carroll2025aioverview; sherrer2025misleading). Similarly, on e-commerce and Q&A websites, individual users typically seek to develop a reputation for expertise and knowledge by providing meaningful answers or reviews and are often rewarded with badges, elite status, or free products in return. [^3] For example, sites like Reddit use a “Karma” system which captures how helpful/unhelpful users found a user’s responses (reddit_karma). Similarly, Yelp recognizes high-quality reviewers by awarding them a “Elite” badge shown publicly on their profile, which in turn gives them access attend exclusive events in the local community as well as other perks (chang2023yelp). However, if consumers stop reading individual reviews/answers, then reviewers/content generators on these sites will lose the opportunity to monetize their content. Thus, the shift to LLM-based summaries threatens the revenue streams of content creators and raises concerns about uncompensated use of their data. Indeed, several major publishers such as The New York Times have started to restrict or revoke AI access to their content, citing concerns that their contributions are being undervalued and misused without fair compensation (nyt_openai_lawsuit_2023). Together, these changes pose challenges to the sustainability of the digital information economy. If content creators are not fairly compensated for their content, then they will have no incentive to provide content to platforms and search engines. This can lead to a general degradation in the quality of AI-powered summaries over time. Thus, while platforms may see short-term benefits from avoiding attribution or compensation to content creators and websites, this is not a sustainable equilibrium in the long term.
@@
As a first step to address this problem, many AI firms have started making licensing deals with large content creators; e.g., recently OpenAI and Murdoch-owned empire of publications like The Wall Street Journal and The New York Post made a deal to allow OpenAI to use the content from these publications (robertson2024openai). Other industry approaches, such as Cloudflare’s recently launched marketplace (cloudflare2025), allow websites to charge AI bots for scraping but determine payments solely by the number of scrape requests rather than the contribution of the scraped content. Overall, there are several challenges associated with such arrangements. First, it is often unclear how to value the body of content from different publishers and arrive at an equitable price for each publisher/contributor. Quantifying a document or publisher’s contribution to a given summary is a non-trivial task – the value that a specific document adds to a summary (in response to a query) depends not only on the document’s own relevance, quality, and reliability but also on the extent to which the information it contains is unique, i.e., the marginal contribution of the document in comparison to other documents. Further, it is also unclear how to aggregate a document’s contributions across queries. For example, one document/publisher may be very valuable for a niche query while another may be moderately valuable for a very popular query. Thus, we need a principled way to correctly value a document’s contribution to both within and across queries. Second, while licensing deals are a potential solution when there are a few large publishers, they are less likely to work when there are a large number of small niche contributors since it is too costly to contract individually with each of them.
@@
Further, for a document valuation approach to be broadly applicable and practical, it needs to satisfy three properties: (1) be summarization procedure agnostic, (2) be evaluation process agnostic, and (3) be scalable and cost-effective. The first property is important because LLM-based summarization procedures vary across platforms (e.g., Google uses Gemini-based systems, whereas Perplexity allows the users to choose their base LLM for summarization). Further summarization techniques and LLMs continue to evolve. Thus, it is essential that any framework we develop is not specific to a single LLM or summarization technique. The second property implies that the framework should be agnostic to the procedure used to score the usefulness of the summary because evaluation methods vary across platforms. For example, search engines typically use implicit feedback, such as how long the user spends on the summary. In contrast, question-answering websites or review aggregators often have explicit feedback on the helpfulness of the summaries. Finally, the third property implies that the document valuation procedure should scale to real platforms and not be unreasonably costly in terms of time or money. Together, these properties help ensure that any document valuation procedure is general and scalable, allowing it to be applied across multiple business contexts. [^4] Note that many simplistic approaches for document valuation can be used, e.g., using the word count of the document, valuation based on a document’s presence/absence in a summary, copyright or link-based attribution, or leave-one-out method suffer from significant disadvantages (ghorbani2019data). However, these methods are unable to properly account for the uniqueness of information, marginal/incremental value, and actual contribution of a document to the summary.
@@
In this paper, we present a framework for equitable document valuation in the context of LLM-based summaries. Our framework addresses the challenges discussed above and satisfies the three properties noted earlier. Our solution concepts builds on the Shapley value framework from cooperative game theory (shapley1953value). First, given a query $q$ and an LLM-based summary based on a specific set of documents, we use the Shapley value framework to distribute the total value generated by the summary (measured by its quality or usefulness) across this set of documents. Shapley values quantify how much each document contributes to the overall summary while accounting for interactions between documents such as redundancy and overlapping information. Formally, Shapley values ensure fair attribution through four properties: Efficiency (the total value is fully distributed across all documents), Symmetry (treat equally contributing documents identically), Null Document (assign zero value to non-contributing documents), and Linearity (additive consistency across multiple evaluations). The first three properties uniquely determine the Shapley value within the context of a given query, while the Linearity property allows us to aggregate Shapley values across queries. As such, this framework addresses both the challenges discussed above. Further, Shapley value calculation is a meta-algorithm, i.e., the exact procedures used for summarization and evaluating the summaries can be black-boxed in Shapley calculations. Thus, it naturally satisfies the first two desirable properties listed earlier.
@@
However, a major hurdle in applying Shapley values to LLM summaries is its exponential computational complexity – if we have $S_{q}$ relevant documents for query $q$, the number of summarizations and evaluations needed to calculate Shapley values is $2^{|S_{q}|}-1$. Since most summaries use anywhere from 4–8 documents for summarization and even small platforms typically process millions of queries, the exact Shapley framework is impractical in real settings. To address this problem, many approximation algorithms have been proposed, such as, Monte Carlo (mann1960values), Truncated Monte Carlo (ghorbani2019data), and Kernel SHAP (lundberg2017unified). However, all these algorithms treat each document as independent entities and do not leverage the textual information of documents. As such, they are unlikely to be efficient.
@@
We propose a novel approximation algorithm – Cluster Shapley – that leverages textual information to improve efficiency while preserving accuracy. The core idea of our approach is intuitive: documents with similar content should have comparable contributions to the final summary and, therefore, should receive similar Shapley values. Instead of treating documents as independent, our method utilizes LLM-generated embeddings to first identify and cluster similar documents. Then, it treats each cluster as a single meta-document for the purposes of the Shapley calculation. Finally, it distributes cluster-level Shapley values equally across all the documents within a cluster. This approach significantly reduces the number of document combinations the LLM must summarize and evaluate without seriously compromising accuracy. Another key advantage of Cluster Shapley is its ability to flexibly trade off computational cost and accuracy via a tunable hyperparameter - clustering diameter $\epsilon$. For tasks requiring high precision, a smaller $\epsilon$ can be selected, while tasks prioritizing speed can utilize a larger $\epsilon$.
@@
We establish a set of theoretical guarantees for our proposed Cluster Shapley algorithm that balance accuracy and computational efficiency. Under a mild Lipschitz continuity assumption in the embedding space, we prove that the approximation error for each document’s Shapley value is bounded by $L\epsilon$, where $L$ is the Lipschitz constant, and the error converges to zero as clusters become finer, i.e., the clustering diameter $\epsilon$ goes to zero. The algorithm reduces the exponential complexity of exact Shapley computation—from $O(2^{n})$ to $O(n^{2}+2^{m})$, where $n=|S_{q}|$ is the number of relevant documents and $m$ is the number of clusters—yielding substantial computational savings when documents exhibit semantic redundancy. Next, we consider settings where $m$ remains large even after the clustering process – usually because the number of relevant documents $n$ is too large to begin with and/or because the set of relevant documents is disparate, and cannot be represented by a small number of clusters. For such settings, we extend our method by incorporating approximation algorithms (e.g., Monte Carlo) at the cluster level. We derive high-probability error bounds that decompose the total approximation error into clustering and cluster-level approximation error. For example, using Monte Carlo sampling for cluster-level Shapley estimation further reduces the total complexity further reduces the computation complexity to $O(n^{2}+m\epsilon^{-2})$ while maintaining the same scale approximation error. Together, these results demonstrate that Cluster Shapley is a principled and scalable solution for equitable document valuation in LLM-generated summaries, enabling efficient deployment even in large-scale settings.
@@
To empirically demonstrate the performance of our algorithm, we apply it to a dataset of products drawn from the popular Amazon product review dataset (hou2024bridging). This dataset closely aligns with real-world applications, as Amazon’s AI-generated review summaries are built on the same repository. We choose 24 products from diverse categories and construct two queries for each product based on frequently mentioned attributes on Amazon, mimicking the information needs of real consumers. For each query, the system summarizes relevant reviews using the RAG framework: the top eight most relevant reviews are retrieved using embedding similarity (between the query and the review), and the LLM generates a query-specific summary based on this retrieved set.
We benchmark our Cluster Shapley algorithm against three widely used approximation methods: Monte Carlo, Truncated Monte Carlo, and Kernel SHAP. Our results show that Cluster Shapley achieves substantial computational savings while maintaining the same accuracy as the benchmark algorithms. Specifically, our method requires only 20–40 unique permutations out of a total of 255 to reach a Mean Absolute Error (MAE) of 0.20, whereas baseline methods — such as Monte Carlo, Truncated Monte Carlo, and Kernel SHAP — require at least 110 permutations to achieve the same accuracy level. Further, if we choose a smaller clustering parameter $\epsilon=0.20$, higher accuracy can be achieved (MAE of 0.0913 and MAPE of 11.85%), while the computation time can still be reduced by 40%. This significant improvement in efficiency suggests that Cluster Shapley can be a practical solution for large-scale platforms.
@@
We also conduct a series of robustness checks and explore extensions to demonstrate the stability and versatility of our algorithm. Our main numerical results remain consistent under a variety of conditions, including alternative evaluation LLMs (e.g., Claude instead of GPT), different error metrics, and different sample splits. We further show that explicitly enforcing intra-cluster similarity leads to improved performance compared to off-the-shelf clustering methods. Finally, we show that our approach easily scales to settings where the platform uses a larger number of documents in the summarization process (large $n$ settings). While standard Shapley is computationally infeasible for large $n$ settings, the version of our Cluster Shapley that incorporates Monte Carlo sampling is not only feasible but is also able to recover sufficiently accurate valuations. Overall, our framework is able to capture the efficiency–accuracy trade-off in Shapley calculations and yields a favorable efficient frontier across a variety of use cases.
In summary, our paper makes three key contributions. First, we address the important and unsolved problem of source document valuation in the context of LLM-generated summaries. By leveraging Shapley values, we propose a framework for fair and transparent document attribution, marking the first study, to the best of our knowledge, that applies this concept to LLM-based search systems. Second, we introduce the Cluster Shapley algorithm, which enhances the efficiency of Shapley value computation by leveraging semantic similarity among documents. We provide theoretical bounds on the performance of this algorithm and demonstrate its empirical performance in a real data setting. We find that the method significantly reduces computational costs while preserving attribution accuracy, making it well-suited for large-scale LLM-driven applications that summarize documents. Finally, the broader idea that Shapley algorithms can be significantly speeded up by considering the similarity between contributors is quite general (extant work treats contributors as independent), and can be applied to other settings, e.g., in machine learning settings, researchers may want to treat observations close to each other in the vector space as similar or belonging to the same cluster, and thereby reduce Shapley computation costs. We thus expect the core idea of using the similarity between contributors to improve Shapley efficiency to be broadly applicable.

### v3

# Fair Document Valuation in LLM Summaries via Shapley Values

@@ Abstract
Large Language Models (LLMs) are increasingly used in systems that retrieve and summarize content from multiple sources, such as search engines and AI assistants. While these systems enhance user experience through coherent summaries, they obscure the individual contributions of original content creators, raising concerns about credit attribution and compensation. We address the challenge of valuing individual documents used in LLM-generated summaries by proposing a Shapley value–based framework for fair document valuation. Although theoretically appealing, exact Shapley value computation is prohibitively expensive at scale. To improve efficiency, we develop Cluster Shapley, a simple approximation algorithm that leverages semantic similarity among documents to reduce computation while maintaining attribution accuracy. Using Amazon product review data, we empirically show that off-the-shelf Shapley approximations, such as Monte Carlo sampling and Kernel SHAP, perform suboptimally in LLM settings, whereas Cluster Shapley substantially improves the efficiency–accuracy frontier. Moreover, simple attribution rules (e.g., equal or relevance-based allocation), though computationally cheap, lead to highly unfair outcomes. Together, our findings highlight the potential of structure-aware Shapley approximations tailored to LLM summarization and offer guidance for platforms seeking scalable and fair content attribution mechanisms.
Keywords: LLMs, Shapley Value, Digital Marketing, Search System, Retrieval Augmented Generation, Revenue Management.
@@ 1 Introduction
The advent of Large Language Models (LLMs) has revolutionized how users search, process, and consume information. Today’s LLM-based search and summarization engines combine the strengths of LLMs with those of traditional search engines. First, unlike traditional search engines that return a list of links in response to a query, LLM-powered search provides a concise summary with references to relevant documents/websites. This frees users from the cognitive load of manually navigating and aggregating information from multiple sources (47). Second, unlike regular LLMs that rely solely on static training data, LLM-based search engines augment their generative process with real-time retrieval and source grounding using a framework known as Retrieval-Augmented Generation (RAG) (17). In the RAG framework, the system first retrieves a set of documents/articles that are most relevant to the query and then uses them as context in the generative process. This ensures that the LLM’s responses are grounded in up-to-date, relevant content and do not suffer from limitations such as non-factual hallucinations and outdated knowledge that plague off-the-shelf LLMs. Thus, LLM-based search and summarization platforms combine the generative aspects of LLMs with the retrieval aspects of search engines by augmenting generation with context/information from documents most relevant to a query.
@@
Over the last few years, all major search engines have adopted these developments and integrated LLMs into their search infrastructures, e.g., Microsoft’s Bing AI (44) and Google’s AI Overview (22). Both these tools provide a summary response to a user’s search query with references or links to relevant websites for informational queries. [^1] Search queries are typically categorized into three groups – (1) informational, where the user is seeking information on a topic, (2) navigational, where the user is simply using the search engine to navigate to a website, and (3) transactional, where the user seeks to perform some transaction, e.g., a purchase. LLM-based summaries are mainly shown for informational queries. Further, new LLM-based search engines such as OpenAI’s ChatGPT Search (47) and Perplexity AI (50) have emerged and grown rapidly in the last few years. The impact of this transformation goes beyond web search and extends to Q&A and informational websites (e.g., Reddit) and e-commerce websites (e.g., Amazon, Best Buy). For example, Amazon now shows an LLM-based summary of all the reviews for a product on the main product page. Customers can further access the source reviews the summary is built on, if needed; see Figure 1 for an example. In addition to broad overviews of the reviews, consumers can query Amazon’s AI-based search tool Rufus for specific information related to the product, and it provides a response based on the information in the reviews/product page (59; 43). Examples of all these applications are illustrated in Web Appendix §A.
@@
The main advantage of LLM-based search and summarization tools is that they simplify the search process for users, who no longer have to click on multiple links and aggregate information from individual sources. This can increase customer satisfaction and lead to higher platform usage (68). For example, (75) empirically shows that LLM-based search and summarization can significantly boost consumer purchases on an e-commerce platform through better articulation of consumer needs and product attributes. Furthermore, these approaches keep customers on the platform’s own website or interface; that is, consumers don’t need to leave the search engine and visit other websites to gather information. This, in turn, can lead to increased time spent on the platform, resulting in more eyeballs and potentially higher revenues (21). For these reasons, the rise of LLM-based search and summarization has generally been seen as beneficial to both platforms and consumers, at least in the short term.
@@
However, the rise of LLM-based automated summarization tools has led to significant concerns among content creators and information providers. In a traditional search model, the source documents (and their corresponding producers) stand to gain traffic and reputation by being shown in response to search queries in organic listings. [^2] Both higher traffic and better reputation can usually be monetized through advertising and/or subscription revenues. In contrast, LLM-powered search delivers a synthesized answer directly in the search interface, allowing users to obtain information without visiting the source websites. For example, after the introduction of Google’s AI overview, website owners have expressed reservations over the loss of traffic and revenues (4; 63; 7). Similarly, on e-commerce and Q&A websites, individual users typically seek to develop a reputation for expertise and knowledge by providing meaningful answers or reviews and are often rewarded with badges, elite status, or free products in return. [^3] For example, sites like Reddit use a “Karma” system which captures how helpful/unhelpful users found a user’s responses (54). Similarly, Yelp recognizes high-quality reviewers by awarding them an “Elite” badge shown publicly on their profile, which in turn gives them access to attend exclusive events in the local community as well as other perks (6). However, if consumers stop reading individual reviews/answers and rating them, the reputation systems on these platforms would cease to function (56), and reviewers/content generators on these sites will lose the opportunity to monetize their content. Thus, the shift to LLM-based summaries threatens the revenue streams of content creators and raises concerns about the uncompensated use of their data. [^4] Indeed, several major publishers such as The New York Times have started to restrict or revoke AI access to their content, citing concerns that their contributions are being undervalued and misused without fair compensation (24). If content creators are not fairly compensated for their content, then they will have no incentive to provide content to platforms and search engines. This can lead to degradation in the quality of AI-powered summaries and failure of platforms’ reputation systems over time. Thus, while platforms may see short-term benefits from avoiding attribution or compensation to content creators and websites, this is not a sustainable equilibrium in the long term.
@@
As a first step to address this problem, many AI firms have started making licensing deals with large content creators; e.g., recently OpenAI and Murdoch-owned empire of publications like The Wall Street Journal and The New York Post made a deal to allow OpenAI to use the content from these publications (57). Other industry approaches, such as Cloudflare’s recently launched marketplace (72), allow websites to charge AI bots for scraping but determine payments solely by the number of scrape requests rather than the contribution of the scraped content. However, there are several challenges associated with such arrangements. First, it is often unclear how to value the body of content from different publishers and arrive at an equitable price for each publisher/contributor. Quantifying a document or publisher’s contribution to a given summary is a non-trivial task – the value that a specific document adds to a summary (in response to a query) depends not only on the document’s own relevance, quality, and reliability but also on the extent to which the information it contains is unique, i.e., the marginal contribution of the document in comparison to other documents. Further, it is also unclear how to aggregate a document’s contributions across queries. For example, one document/publisher may be very valuable for a niche query while another may be moderately valuable for a very popular query. Thus, we need a principled way to correctly value a document’s contribution to both within and across queries. Second, while licensing deals are a potential solution when there are a few large publishers, they are less likely to work when there are a large number of small niche contributors since it is too costly to contract individually with each of them.
@@
Further, for a document valuation approach to be broadly applicable and practical, it needs to satisfy three properties: (1) be summarization procedure agnostic, (2) be evaluation process agnostic, and (3) be scalable and cost-effective. The first property is important because LLM-based summarization procedures vary across platforms (e.g., Google uses Gemini-based systems, whereas Perplexity allows the users to choose their base LLM for summarization). Further summarization techniques and LLMs continue to evolve. Thus, it is essential that any framework we develop is not specific to a single LLM or summarization technique. The second property implies that the framework should be agnostic to the procedure used to score the usefulness of the summary because evaluation methods vary across platforms. For example, search engines typically use implicit feedback, such as how long the user spends on the summary. In contrast, question-answering websites or review aggregators often have explicit feedback on the helpfulness of the summaries. Finally, the third property implies that the document valuation procedure should scale to real platforms and not be unreasonably costly in terms of time or money. Together, these properties help ensure that any document valuation procedure is general and scalable, allowing it to be applied across multiple business contexts. [^5] Note that many simplistic approaches for document valuation can be used, e.g., using the word count of the document, valuation based on a document’s presence/absence in a summary, copyright or link-based attribution, or leave-one-out method suffer from significant disadvantages (20). However, these methods are unable to properly account for the uniqueness of information, marginal/incremental value, and actual contribution of a document to the summary.
@@
In this paper, we present a framework for fair document valuation in the context of LLM summaries. Our framework addresses the challenges discussed above and satisfies the three properties noted earlier. Our solution concept builds on the Shapley value framework from cooperative game theory (62). First, given a query $q$ and an LLM summary based on a specific set of documents, we use the Shapley value framework to distribute the total value generated by the summary (measured by its quality or usefulness) across this set of documents. Shapley values quantify how much each document contributes to the overall summary while accounting for interactions between documents, such as redundancy and overlapping information. Formally, Shapley values ensure fair attribution through four properties: Efficiency (the total value is fully distributed across all documents), Symmetry (treat equally contributing documents identically), Null Document (assign zero value to non-contributing documents), and Linearity (additive consistency across multiple evaluations). Note that this framework supports value attribution at the query level, and the linearity property enables aggregation of values across queries. This flexibility is crucial for revenue attribution, as it accommodates both per-query revenues (e.g., advertising or API calls) and aggregated subscription-based revenues. As such, this framework addresses both the challenges discussed above. Moreover, because Shapley value computation treats the summarization and evaluation procedures as black boxes, it remains compatible with any LLM summarization method or evaluation metric, satisfying the first two design requirements introduced earlier.
@@
However, a major hurdle in applying Shapley values to LLM summaries is its exponential computational complexity – if we have $S_{q}$ relevant documents for query $q$, the number of summarizations and evaluations needed to calculate Shapley values is $2^{|S_{q}|}-1$. Even for modest-scale platforms that process millions of queries, this becomes computationally infeasible in practice. To address this problem, many approximation algorithms have been proposed, such as Monte Carlo (42), Truncated Monte Carlo (20), and Kernel SHAP (40). However, all these algorithms treat each document as independent entities and do not leverage the textual information of documents. As a result, their empirical performance in LLM summarization settings is often suboptimal.
@@
In this paper, we develop a simple yet effective approximation algorithm, Cluster Shapley, that leverages textual information to improve computational efficiency. The intuition is straightforward: documents with similar content are likely to make comparable contributions to the final summary and therefore should receive similar Shapley values. Instead of treating documents as independent, the algorithm uses LLM-generated embeddings to group semantically similar documents into clusters. Each cluster is then treated as a single meta-document in the Shapley calculation, and the resulting cluster-level values are distributed evenly across the documents within each cluster. This reduces the number of combinations the LLM must summarize and evaluate, providing efficiency gains with limited loss in accuracy. A tunable hyperparameter-the clustering diameter $\epsilon$-controls the balance between accuracy and cost: smaller values of $\epsilon$ yield more accurate but slower computations, whereas larger values offer faster, coarser approximations.
@@
We establish a set of theoretical guarantees for our proposed Cluster Shapley algorithm that balance accuracy and computational efficiency. Under a mild Lipschitz continuity assumption in the embedding space, we prove that the approximation error for each document’s Shapley value is bounded by $L\epsilon$, where $L$ is the Lipschitz constant, and the error converges to zero as clusters become finer, i.e., the clustering diameter $\epsilon$ goes to zero. The algorithm reduces the exponential complexity of exact Shapley computation-from $O(2^{n})$ to $O(n^{2}+2^{m})$, where $n=|S_{q}|$ is the number of relevant documents and $m$ is the number of clusters-yielding substantial computational savings when documents exhibit semantic redundancy. Next, we consider settings where $m$ remains large even after the clustering process – usually because the number of relevant documents $n$ is too large to begin with and/or because the set of relevant documents is disparate, and cannot be represented by a small number of clusters. For such settings, we extend our method by incorporating approximation algorithms (e.g., Monte Carlo) at the cluster level. We derive high-probability error bounds that decompose the total approximation error into clustering and cluster-level approximation error. For example, using Monte Carlo sampling for cluster-level Shapley estimation further reduces the total complexity further reduces the computation complexity to $O(n^{2}+m\epsilon^{-2})$ while maintaining the same-order approximation error. Together, these results demonstrate that Cluster Shapley is a principled and scalable solution for fair document valuation in LLM-generated summaries, enabling efficient deployment even in large-scale settings.
@@
To empirically demonstrate the performance of our algorithm, we apply it to a dataset of products drawn from the popular Amazon product review dataset (29). This dataset closely aligns with real-world applications, as Amazon’s AI-generated review summaries are built on the same repository. We choose 24 products from diverse categories and construct two queries for each product based on frequently mentioned attributes on Amazon, mimicking the information needs of real consumers. For each query, the system summarizes relevant reviews using the RAG framework: the top eight most relevant reviews are retrieved using embedding similarity (between the query and the review), and the LLM generates a query-specific summary based on this retrieved set.
We benchmark the proposed Cluster Shapley algorithm against two simple attribution rules-equal attribution and relevance-weighted attribution-and three widely used approximation methods: Monte Carlo (42), Truncated Monte Carlo (20), and Kernel SHAP (40). The results yield two main insights. (1) Off-the-shelf approximation algorithms are empirically far from optimal in the context of LLM-based summarization. These algorithms, originally designed for settings with independent contributors, fail to exploit the rich semantic structure of textual data and therefore deliver poor efficiency–accuracy trade-offs. In contrast, Cluster Shapley substantially improves the efficient frontier: for instance, our method requires only 20–40 unique permutations (out of 255 total) to reach an MAE of 0.20, whereas other approximation methods need more than 110 permutations to achieve comparable accuracy. This finding underscores the importance of developing Shapley approximation algorithms specifically tailored for LLM summarization-leveraging embedding-based similarity and other contextual structures-and points to a promising direction for future research. (2) Simple attribution rules are highly biased, raising fairness concerns in practical applications. Equal attribution, despite its negligible computational cost, results in an MAPE exceeding 150%, and even relevance-weighted attribution remains above 30%. These large deviations from the fair Shapley baseline demonstrate that simplistic heuristics, though computationally cheap, can substantially distort individual document contributions. Together, these findings highlight that more effective, structure-aware Shapley approximations are both feasible and necessary: they enable fairer value attribution with affordable computational cost.
@@
We also conduct a series of robustness checks and explore extensions to demonstrate the stability and versatility of our algorithm. Our main numerical results remain consistent under a variety of conditions, including alternative evaluation LLMs (e.g., Claude instead of GPT), different error metrics, and different sample splits. We further show that explicitly enforcing intra-cluster similarity leads to improved performance compared to off-the-shelf clustering methods. Finally, we show that our approach easily scales to settings where the platform uses a larger number of documents in the summarization. While standard Shapley is computationally infeasible for large $n$ settings, the version of our Cluster Shapley that incorporates Monte Carlo sampling is not only feasible but is also able to recover sufficiently accurate valuations. Overall, our framework is able to capture the efficiency–accuracy trade-off in Shapley calculations and yields a favorable efficient frontier across a variety of use cases.
In summary, our paper makes four key contributions. First, we address the important and largely unsolved problem of source document valuation in the context of LLM-generated summaries. By leveraging the Shapley value framework, we propose a principled approach for fair document valuation, to the best of our knowledge, the first study to apply this concept to LLM-based summarization systems. Second, we introduce the Cluster Shapley algorithm, which improves the computational efficiency of Shapley value approximation by exploiting semantic similarity among documents. We provide theoretical guarantees on its performance and evaluate its empirical effectiveness using real-world data, showing that it substantially improves the efficiency–accuracy frontier. Third, our results open up a new research direction by demonstrating that Shapley approximation algorithms can be significantly improved in LLM applications when additional structure, such as embedding-based similarity, is available. This highlights the potential for developing a broader class of structure-aware Shapley approximation algorithms for LLM applications. Fourth, we show that simple attribution rules, though computationally cheap, can lead to substantial unfair outcomes. This insight provides practical guidance for decision-makers: platforms relying on simplistic heuristics may systematically undervalue or overvalue individual contributors, underscoring the importance of adopting fair and theoretically grounded attribution frameworks.